OpenStudy (anonymous):

please help Find all the real fourth roots of 625

OpenStudy (anonymous):

for real? lol one of them should come to mind... considering \[5^4 = 625\] ^.^

OpenStudy (anonymous):

how do i solve it? you dont have to give me the answer id like to learn.

OpenStudy (anonymous):

Getting fourth roots? No method exists that I know of... except knowing it by heart... ^.^

OpenStudy (anonymous):

oops real 4th roots, sorry factor maybe

OpenStudy (anonymous):

Or you could get the square root twice... that's the same.. :D

OpenStudy (anonymous):

what are cube roots?

OpenStudy (anonymous):

r is the cube root of a number x if r^3 = x

OpenStudy (anonymous):

finding the cube roots of 1000/27

OpenStudy (anonymous):

for instance, 2 is the cube root of 8. because 2^3 = 8

OpenStudy (anonymous):

so id have to find simplify the fraction then solve

OpenStudy (anonymous):

Yes, actually, factor the numerator and denominator...and then to get the cube root, divide all exponents by 3...

OpenStudy (anonymous):

i got 37

OpenStudy (anonymous):

Not possible... 37^3 = 50653

OpenStudy (anonymous):

come on, what is 1000 in factored form?

OpenStudy (anonymous):

1/1000

OpenStudy (anonymous):

1000/1

OpenStudy (anonymous):

huh? lol well, this isn't exactly fully factored, but 1000 = 10^3 yeah?

OpenStudy (anonymous):

yeah but they gave me a fraction of 1000/27

OpenStudy (anonymous):

hold your horses :P the denominator 27...could you factor that?

OpenStudy (anonymous):

7^3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

7^3 is 343 :P

OpenStudy (anonymous):

3^3

OpenStudy (anonymous):

brain fart

OpenStudy (anonymous):

yeah!\[\large \frac{1000}{27}=\frac{10^3}{3^3}\] So, what is the cube root now? ^.^

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

To get the cube root, just divide ALL exponents by 3.

OpenStudy (anonymous):

3.3333/1?

OpenStudy (anonymous):

I said the EXPONENTS, not the bases :P

OpenStudy (anonymous):

ohhhh 1

OpenStudy (anonymous):

so its 10/3

OpenStudy (anonymous):

That's right ^.^

OpenStudy (anonymous):

wow i feel dumb.

OpenStudy (anonymous):

never really paid attention in math class. now its paying me back.

OpenStudy (anonymous):

lol me neither... math class is boooring ^.^

OpenStudy (anonymous):

but thank you. sorry i made it so difficult for you.

OpenStudy (anonymous):

lol don't worry about it ^.^

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

What about your original question, then? the real fourth roots of 625?

OpenStudy (anonymous):

yes but i got the both now. thank you

OpenStudy (anonymous):

20 more to go. lol

OpenStudy (anonymous):

i hate algebra

OpenStudy (anonymous):

I suppose it hates you too :)

OpenStudy (anonymous):

Duh.